OpenStudy (anonymous):

For the equation x2 + 3x + j = 0, find all the values of j such that the equation has two real number solutions. Can someone please show me the steps of how to do this? Thanks! :)

Parth (parthkohli):

I think that factoring is the best solution, and if you think about it, then we can put j as 2. 2 * 1 = 2 2 + 1 = 3 j = 2

Parth (parthkohli):

Let's try to factor it further. \(\Large \color{Black}{\Rightarrow x^2 + 2x + x + 2 }\) \(\Large \color{Black}{\Rightarrow x(x + 2) + 1(x + 2) }\) \(\Large \color{Black}{\Rightarrow (x + 1)(x + 2) }\) x = -1 or -2(two solutions :D)

OpenStudy (anonymous):

Thanks so much!